Thread View: sci.astro.hubble
1 messages
1 total messages
Started by hst-status
Mon, 28 Dec 1992 18:18
HST STATUS REPORT, 12/22 - 12/23
Author: hst-status
Date: Mon, 28 Dec 1992 18:18
Date: Mon, 28 Dec 1992 18:18
161 lines
6371 bytes
6371 bytes
HUBBLE SPACE TELESCOPE
DAILY REPORT #790
PERIOD COVERED: 0800 EST 12/22/92 - 0800 EST 12/23/92
Daily Status Report as of 358/1200Z.
1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED:
1.1 Completed GHRS/2 2536 (Deuterium in the Local Interstellar Gas)
The High Resolution Spectrograph used its Echelle spectrograph
to make observations of interstellar absorption lines in the spectra
of white dwarf star G191-B2B. The first four observations were: two
deflection calibrations, a spiral search, and the peakup. There were 15
spectral observations which followed. All of the observations were
made in course track, and all were reported to be okay.
1.2 Completed Two Sets of PC 4714 (Internal Delta Flats)
A total of nine internal flat observations were made with the
Planetary Camera. These images will be used for the restoration of
Earth flats, which are altered due to the changes in quantum
efficiency which occur after a WFPC decontamination operations. All
of the internal flats were received and displayed. They were all
okay.
1.3 Completed Two Sets of S/C 4215 (FGS ITS Test)
The Fine Guidance Sensor internal test was executed as scheduled,
and no problems were reported.
1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871
Uplinked at 358/0100Z.
1.5 Completed FOS/RD 3816 (White Dwarf Stars)
The Faint object Spectrograph was used to obtain spectral
observations of white dwarf star L151-81. The onboard target
acquisition was successful. The OSS monitored the first spectral
observation, and it was successful. The other observations occurred
during LOS, and still are on the science tape recorder.
1.6 Completed PC 4106 (HST Medium Deep Survey: Low Latitude)
The Planetary Camera was used to make low latitude observations in
parallel with the prime science observations. This is to provide for
serendipitous discoveries and data for a basic sky survey. The
observations were monitored in the OSS, and no problems were
reported.
1.7 Completed WFC-PC 4109 (Rapid Internal Monitor)
Both the Wide Field Camera and the Planetary Camera were used to
obtain internal flat observations using filter F555W. This operation
is to check for different forms of contamination (i.e., daises and
measles). The internal flat observations were monitored in the OSS,
and no problems were reported.
2.0 FLIGHT OPERATIONS SUMMARY:
2.1 Guide Star Acquisitions:
Scheduled Acquisitions : 1
Successful : 1
Scheduled Reacquisitions : 7
Successful : 7
Losses of Lock : 0
2.2 FHST Updates:
Scheduled : 7
Successful : 7
2.3 Significant HSTARs: None
2.4 Operations Note: None
3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD:
3.1 Complete Two Sets of GHRS 2536 (Deuterium in the Local Interstellar
Gas)
3.2 Complete Two Sets of PC 3912 (Cores of Early Type Galaxies)
3.3 Complete FOS/BL 3939 (Lyman Alpha Regions in QSOs)
3.4 Complete HSP/UV2 1099 (Active Galactic Nuclei)
3.5 Complete Uplink Updated Ephemeris Tables, Per MOSES 2871
3.6 Complete Two Sets of FGS 3688 (Diagnostic of Early Type Stellar
Evolution)
3.7 Complete HSP/UV1-PC 4741 (Orbital Effects Test Calibration)
3.8 Complete One Set of FOS/RD 3816 (White Dwarf Stars) and Initiate a
Second.
3.9 Complete PC 4106 (HST Medium Deep Survey: Low Latitude)
3.10 Complete PC 4347 (Flood Retention Test)
3.11 Complete FOS/RD 3566 (Do Galaxies Produce Absorption Lines)
3.12 Complete WFC 4105 (HST Medium Deep Survey: High Latitude)
3.13 Complete PC 4712 (Measles Monitor)
3.14 Complete HSP/UV1 3257 (Periodic Variations in DQ Herculis Stars)
3.15 Complete PC 3657 (The Host Galaxies of the BL Lacertea Objects)
3.16 Complete PC 4349 (Black Holes, Stellar Dynamics and Populations)
3.17 Complete PC 4140 (Rapid Internal Monitor - Part 2)
3.18 Complete PC 3286 (Nuclei of Nearly Normal Galaxies)
3.19 Complete FOC/96 4107 (Search for QSOs Suitable for Subsequent
Observations)
3.20 Complete WFC-PC 4109 (Rapid Internal Monitor)
3.21 Complete S/C 4215 (FGS ITS Test)
3.22 Complete HRS/2 4065 (Dark Noise Monitoring)
3.23 Complete FOC/96 3504 (Resolution Imaging and Spectroscopy of AGN)
3.24 Complete WFC 2227 (Determination of Extra-galactic Distance Scale:
M-81)
3.25 Complete GHRS/2 3706 (Nature of Gaseous Loops in the Milky Way
Halo)
4.0 ANOMALY STATUS:
4.1 Significant New Anomalies: None
4.2 Significant Unresolved Anomalies: None
5.0 SIGNIFICANT FORTHCOMING EVENTS:
There will be a partial eclipse of the Sun at the end of Dec. 23
(today) and the start of Dec. 24 (tomorrow). This will effect two
orbital passes of the Space Telescope, one orbit in each day. The
First orbit will be effected for a 20 minute period, with a maximum
of 37% reduction in Sun light. The second orbit will be effected for
six minutes, with a 2.4% reduction in Sun light. The Battery Rate of
Charge Test will be disabled in the EPS System, and the Course Sun
Sensors in the PCS System.
Due to the Christmas and New Years holidays no daily reports will be
written on Thursday Dec. 24, Friday Dec. 25, and Friday Jan. 1. A
daily report will be written Monday Through Thursday (Dec. 28 - Dec.
31).
Thread Navigation
This is a paginated view of messages in the thread with full content displayed inline.
Messages are displayed in chronological order, with the original post highlighted in green.
Use pagination controls to navigate through all messages in large threads.
Back to All Threads